Easy problem. This is a timed assignment 30mins answer with work will be appreciated

15. If the magnitade of is 10, and the magnitude of scalar product of the two vectors shown is 20, und a 120°, determine the a. +100 b.-100 c. +8 e.-35
Answer #1

A = 10 i

B = 20 at 120 deg from -ve x-axis = 20 st 60 deg from +ve x-axis

B = 20*cos 60 deg i + 20*sin 60 deg

B = 20*(1/2) i + 20*(sqrt 3)/2 j

B = 10 i + 10*sqrt 3 j

A.B = (10 i).(10 i + 10*sqrt 3 j)

we know that in scalar product

i.i = 1 and i.j = 0

A.B = 10*10*(i.i) + 0 = +100

Correct option is A.